# Towards a constructive framework for control theory

### v1

@@ I Introduction
[…]
@@ III Optimal control and stabilization
[…]
@@ IV Danskin’s theorem
[…]
@@
For the continuity part, let $\mu^{x}_{\varphi}$ be the continuity modulus of $\varphi$ w. r. t. $x$ to mean: $\forall\theta\in\Theta\,\,\varphi(x,\theta)-\varphi(y,\theta)\leq\mu^{x}_{\varphi}(\left\lVert x-y\right\rVert)$. In particular, the latter holds with $\theta^{\delta}(x)$, a $\delta$-optimizer for an arbitrary $\delta>0$, in place of $\theta$. Observe that $\varphi(x,\theta^{\delta}(x))-\varphi(y,\theta^{\delta}(x))\geq\varphi(x,\theta^{\delta}(x))-\max_{\theta\in\Theta}\varphi(y,\theta)$. So, $\varphi(x,\theta^{\delta}(x))-\max_{\theta\in\Theta}\varphi(y,\theta)\leq\mu^{x}_{\varphi}(\left\lVert x-y\right\rVert)$. Now, using a $\delta$-optimizer at $y$, we have: $\varphi(x,\theta^{\delta}(x))-\varphi(y,\theta^{\delta}(y))\leq\mu^{x}_{\varphi}(\left\lVert x-y\right\rVert)+\delta$. Thus, $\psi(x)-\psi(y)\leq\mu^{x}_{\varphi}(\left\lVert x-y\right\rVert)+\delta+2\delta$, where the last two $\delta$s relate $\psi(x),\psi(y)$ to their respective $\delta$-maximal values. Reversing the order of $x,y$ and observing that $\delta$ was arbitrary, it follows that $\left\lvert\psi(x)-\psi(y)\right\rvert\leq\mu^{x}_{\varphi}(\left\lVert x-y\right\rVert)$ as required. Fix an $x,v,\delta>0$ and observe the following:
@@
Since $\varphi$ is continuously differentiable, for any $\varepsilon_{1}>0$ there is an ¯$\bar{}\varepsilon>0$ s. t.
@@
[…]
@@ V Selector theorems
Recently, we showed that extraction of measurable selectors could also be made constructive [40]. Whereas the full details can be found in the related work, let us outline the result in this section. First, fix a compact $\mathbb{X}\in\mathbb{R}^{n}$ and let a block be a not necessarily non-empty (closed) hyperrectangle with rational vertices in $\mathbb{R}^{n}$. Let unions of blocks be called generalized blocks. For a generalized block $\mathbb{B}=\bigcup_{i}\mathbb{B}_{i}$, if any block $A\in\mathbb{B}$ is inside finitely many $\mathbb{B}_{i}$s, $\mathbb{B}$ is called locally finitely enumerable (or just finitely enumerable, if $\mathbb{B}$ is finite as a sequence). Notice, when dealing with $\mathbb{B}=\bigcup_{i}\mathbb{B}_{i}$, we always assume the underlying sequence $\{\mathbb{B}_{i}\}_{i}$ be available. If $\mathbb{B}$ is locally finitely enumerable and $\{\mathbb{B}_{i}\}_{i}$ are disjoint, the generalized block is called proper. For a generalized block $\mathbb{B}=\bigcup_{i}\mathbb{B}_{i}$ define a map $\mu(\mathbb{B})\triangleq\sum_{i}\mu(\mathbb{B}_{i})$ where $\mu(\mathbb{B}_{i})$ is the volume of the respective hyperrectangle (notice the map $\mu$ generalizes the definition [27, Chapter 6], and thus is not treated as the classical Lebesque measure).
We say a sequence of generalized blocks =$\mathcal{E}=\{\mathbb{B}^{j}\}_{j}$ is a representable domain in $\mathbb{X}$ if for any $\varepsilon>0$, there exists another generalized block $\mathbb{J}$ with $\mu(\mathbb{J})\leq\varepsilon$ s. t. $\partial_{\mathcal{E}}\Subset\mathbb{J}\setminus\partial_{\mathcal{\mathbb{J}}}$ and $\mathbb{X}\setminus\mathbb{J}\subseteq{\mathcal{E}}$, where $\partial_{\partial E}$ is the generalized block which is the union of the boundaries of all $\mathbb{B}^{j}_{i}$s and $A\Subset B$ means (constructively) well-contained , i. e., $\exists\lambda>0\,\,A+\lambda\subseteq B$. We will consider in the following measurable (single- and set-valued) functions whose domains are proper and representable. The idea behind representable domains is that they entail an algorithms which give an “arbitrarily thin” generalized blocks s. t. these domains cover the total space minus these generalized blocks.
@@ VI System analysis
Let $\mathbb{X}\subset\mathbb{R}^{n}$ be compact and $\dot{x}=f(x,t),x\in\mathbb{X}$ be a dynamical system with the equilibrium point $x_{e}=0$ and $f$ Lipschitz-continuous in $x$. Suppose there exist positive-definite functions $V,w_{1},w_{2},w_{3}:\mathbb{X}\times\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}$, $V$ continuously differentiable, $w_{1},w_{2},w_{3}$ strictly increasing, $w_{3}$ Lipschitz continuous, with the following properties:

### v2

@@ I Introduction
Notation and abbreviations. Convergence of $t$ to $a$ from the right: $t\searrow a$. A closed ball with a radius $r$ centered at $x$: $\mathcal{B}_{r}(x)$, or just $\mathcal{B}_{r}$ if $x=0$. A closed hypercube with a side length $r$ centered at $x$: $\mathcal{H}_{r}(x)$, or just $\mathcal{H}_{r}$ if $x=0$. Euclidean distance, or $\sup$-norm of a function: $\left\lVert\bullet\right\rVert$. Domain of a function: dom. “Such that”: s. t. . “With respect to”: w. r. t.. “Without loss of generality”: w. l. g..
@@ III Optimal control and stabilization
The justification of Theorem 1 is that physically every signal is bounded and has a finite rate of change. For a generic optimal control problem $\min_{\kappa\in\mathcal{U}}J[\kappa]\,\,\text{s.\,t.\,}\mathcal{D}x=f(x,u)$, where $\mathcal{D}$ is a suitable differential operator, we can apply Theorem 1 if $\mathcal{U}$ is a located subset of a space of all equi-Lipschitz and equi-bounded functions (with the aid of [27, Lemma 4.3]). Otherwise, if the system is , e. g., input-to-state stable, we have $\forall t\,\,\forall\kappa\in\mathcal{U}\,\,\left\lVert x(t)\right\rVert\leq\alpha(\left\lVert x(0),t\right\rVert)+\beta(\left\lVert\kappa\right\rVert)$ with $\alpha$ of class $\mathcal{KL}$ and $\beta$ – of $\mathcal{K}$. In this case, we may derive a uniform bound on $\left\lVert x\right\rVert$ for all policies in $\mathcal{U}$, using their common bound, and apply Theorem 1 directly. Further examples of applications of this constructive theorem include dynamic programming and reinforcement learning and may be found in [26].
@@ IV Danskin’s theorem
We will use the following directional super-derivative (of a function $\psi:\mathbb{R}^{n}\rightarrow\mathbb{R}$ in direction of a vector $v$) will be used: $\mathcal{D}^{+}_{v}\psi(x)\triangleq\limsup_{\varepsilon\searrow 0}\tfrac{\psi(x+\varepsilon v)-\psi(x)}{\varepsilon}$. By analogy, $\liminf$ in the above will be used for the directional sub-derivative $\mathcal{D}^{-}_{v}\psi(x)$. If both coincide, the common limit is simply the directional derivative $\mathcal{D}_{v}\psi(x)$.
@@
[…]
@@
Since $\varphi$ is continuously differentiable, for any $\varepsilon_{1}>0$ there is an $\bar{}\varepsilon>0$ s. t.
@@
But since $\theta^{\varepsilon\delta}(x)$ was arbitrary from $E^{\delta}_{\Theta}(x)$ (provided that $\varepsilon<1$) we may write $\max_{\theta\in E^{\delta}_{\Theta}(x)}$ in front of the quotient in the right-hand side of the above. The rest of the argument is the same as before, but applying $\liminf$ instead of $\limsup$ and noticing that $\delta$ was arbitrary, yielding $\mathcal{D}^{-}_{v}\psi(x)\geq\max_{\theta\in E^{\delta}_{\Theta}(x)}\mathcal{D}_{v}\varphi(x,\theta)$. Combining this with (11) gives the result. ∎
@@ V Selector theorems
Recently, we showed that extraction of measurable selectors could also be made constructive [40]. Whereas the full details can be found in the related work, let us outline the result in this section. Let a block be a not necessarily non-empty (closed) hyperrectangle with rational vertices in $\mathbb{R}^{n}$. Let unions of blocks be called generalized blocks. For a generalized block $\mathbb{B}=\bigcup_{i}\mathbb{B}_{i}$, if any block $A\in\mathbb{R}^{n}$ intersects with only finitely many $\mathbb{B}_{i}$s, $\mathbb{B}$ is called locally finitely enumerable (or just finitely enumerable, if $\mathbb{B}$ is finite as a sequence). Notice, when dealing with $\mathbb{B}=\bigcup_{i}\mathbb{B}_{i}$, we always assume the underlying sequence $\{\mathbb{B}_{i}\}_{i}$ be available. If $\mathbb{B}$ is locally finitely enumerable and $\{\mathbb{B}_{i}\}_{i}$ are disjoint, the generalized block is called proper. Let $\mathbb{B}=\bigcup_{i}\mathbb{B}_{i}$ be a locally finitely enumerable generalized block, and $A$ be a block. Define a map $\mu_{A}(\mathbb{B})\triangleq\sum_{i}\mu(\mathbb{B}_{i}\cap A)$ where $\mu(\mathbb{B}_{i}\cap A)$ is the volume of the respective hyperrectangle (possibly empty) resulting from the intersection. Notice the map $\mu$ generalizes the definition [27, Chapter 6], and thus is not treated as the classical Lebesgue measure.
Fix a compact set $\mathbb{X}\subset\mathbb{R}^{n}$. Hence, there exists a block $A$ which contains $\mathbb{X}$. If a generalized block $\mathbb{B}=\{\mathbb{B}_{i}\}_{i}$ is such that each $B_{i}$ is inside $\mathbb{X}$, we simply write $\mu(\mathbb{B})$ meaning a suitable ambient bounding block $A$.
We say a sequence of generalized blocks $\mathcal{E}=\{\mathbb{B}^{j}\}_{j}$ is a representable domain in $\mathbb{X}$ if for any $\varepsilon>0$, there exists another generalized block $\mathbb{J}\in\mathbb{X}$ with $\mu(\mathbb{J})\leq\varepsilon$ s. t. $\partial_{\mathcal{E}}\Subset\mathbb{J}\setminus\partial_{\mathcal{\mathbb{J}}}$, and $\mathbb{X}\setminus\mathbb{J}$ exists and satisfies $\mathbb{X}\setminus\mathbb{J}\subseteq{\mathcal{E}}$, where $\partial_{\mathcal{E}}$ is the generalized block which is the union of the boundaries of all $\mathbb{B}^{j}_{i}$s and $A\Subset B$ means (constructively) well-contained , i. e., $\exists\lambda>0\,\,A+\lambda\subseteq B$.
We will consider in the following measurable (single- and set-valued) functions whose domains are proper, representable, and whose values are located.
The idea behind representable domains is that they entail an algorithm which yields “arbitrarily thin” generalized blocks in a way that these domains cover the total space minus the said generalized blocks. Noticing $F(x)$ are located, let us introduce the following definition.
@@ VI System analysis
Let $\mathbb{X}\subset\mathbb{R}^{n}$ be compact and $\dot{x}=f(x,t),x\in\mathbb{X}$ be a dynamical system with the equilibrium point $x_{e}=0$ and $f$ Lipschitz-continuous in $x$. Suppose there exist positive-definite functions $w_{1},w_{2},w_{3}:\mathbb{X}\times\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}, VV$ continuously differentiable, $w_{1},w_{2},w_{3}$ strictly increasing, $w_{3}$ Lipschitz continuous, with the following properties: